# Totally geodesic submanifolds of the homogeneous nearly Kähler 6-manifolds and their $\mathsf{G}_{2}$-cones

### v2

@@ 6. Riemannian cones and totally geodesic submanifolds
In this section we start by recalling the definition and the basic properties of Riemannian cones. After that we will prove a structure result for totally geodesic submanifolds of Riemannian cones. We refer the reader to [42] for a detailed account on semi-Riemannian cones.
Let $M$ be a Riemannian manifold, which for our purposes will be assumed to be real analytic and complete. We define its Riemannian cone as the warped product $\widehat{M}=\mathbb{R}^{+}\times_{f}M$, where $f\colon\mathbb{R}^{+}\to\mathbb{R}^{+}$ is the identity map. More explicitly, if $g=\langle\cdot,\cdot\rangle$ denotes the metric on $M$ and $r\colon\widehat{M}\to\mathbb{R}^{+}$ is the projection on the first factor, the inner product on $\widehat{M}$ is given by $\widehat{g}=\mathrm{d}r^{2}+r^{2}g$.
@@
We can also describe the geodesics of $\widehat{M}$ in terms of those of $M$. Let $(\tau,p)\in\widehat{M}$ be any point and consider the tangent vector $w=a\partial_{r}+v$, where $a\in\mathbb{R}$ and $v\in T_{p}M$ are arbitrary. From [42, Equation 2.7], we see that the geodesic $\widehat{\gamma}(t)=\widehat{\exp}_{(\tau,p)}(tw)=(\rho(t),\alpha(t))$ is given in a neighbourhood of $t=0$ by
@@
Now, let $f\colon\widehat{M}\to\widehat{N}$ be an isometry and fix $(\tau,p)\in\widehat{M}$ with image $(s,q)=f(\tau,p)$. Since $f$ is an isometry, it sends $\mathcal{C}_{(\tau,p)}$ to $\mathcal{C}_{(s,q)}$, and thus we have $f_{*(\tau,p)}((\partial_{r})_{(\tau,p)})=\pm(\partial_{r})_{(s,q)}$. The equality $f_{*(\tau,p)}((\partial_{r})_{(\tau,p)})=~{}-(\partial_{r})_{(s,q)}$ is not possible, because in that case the maximal geodesic $\gamma(t)=\widehat{\operatorname{exp}}_{(\tau,p)}(t\partial_{r})$ would be mapped to the maximal geodesic $\beta(t)=\widehat{\operatorname{exp}}_{(s,q)}(-t\partial_{r})$, which is not possible because the first geodesic is defined for all $t\geq 0$, while the second one is not. We deduce that $f_{*(\tau,p)}((\partial_{r})_{(\tau,p)})=~{}(\partial_{r})_{f(\tau,p)}$. Plus, the maximal geodesic $\gamma(t)=\widehat{\operatorname{exp}}_{(\tau,p)}(t\partial_{r})$ is mapped to the maximal geodesic $\beta(t)=~{}\widehat{\operatorname{exp}}_{(s,q)}(t\partial_{r})$, and their corresponding intervals of definition are $(-\tau,0)$ and $(-s,0)$. Hence, $\tau=s$. All in all, we have seen that $f$ sends the link $\{\tau\}\times M$ to $\{\tau\}\times N$ for each $\tau\in\mathbb{R}^{+}$, so $f$ takes the form $f(\tau,p)=(\tau,h(\tau,p))$ for a map $h\colon\widehat{M}\to N$. Furthermore, at each $(\tau,p)\in\widehat{M}$ we have
@@
Many geometric properties of Riemannian manifolds can be translated into geometric properties of their cones. For instance, as a consequence of (9), $M$ is an Einstein manifold with $\operatorname{Ric}=(n-1)\langle\cdot,\cdot\rangle$ if and only if $\widehat{M}$ is Ricci-flat. It turns out that nearly Kähler structures on six-dimensional manifolds are related to $\mathsf{G}_{2}$-structures on their cones. We briefly describe this relationship, see [5] for details. For the general theory of $\mathsf{G}_{2}$-manifolds, we refer the reader to [33, Chapter 10].
@@
There is also a relationship between submanifolds of $M$ that have a nice interaction with $J$ and calibrated cones inside $\widehat{M}$. The notions of calibrated geometry were introduced in the seminal paper [27] by Harvey and Lawson. We remind that a calibration on a Riemannian manifold $N$ is a closed differential form $\omega\in\Omega^{k}(M)$ satisfying $\omega(v_{1},\dots,v_{k})\leq 1$ whenever $v_{1},\dots,v_{k}$ are unit vectors in $TN$. A $k$-dimensional oriented submanifold $S$ of $N$ is calibrated if the restriction of $\omega$ to $S$ is equal to the Riemannian volume form of $S$, and it follows that $S$ is a minimal submanifold. It can be shown that for the case of a $\mathsf{G}_{2}$-manifold $(N,\phi)$, both $\phi$ and its Hodge dual $\star\phi$ are calibrations ([27, Theorem 1.4 and Theorem 1.16]). We say in this case that $S$ is associative (respectively, coassociative) if it is calibrated with respect to $\phi$ (respectively, $\star\phi$). Coming back to the case that $N=\widehat{M}$ is the cone of a six-dimensional strict nearly Kähler manifold, it is known that the cone of a $J$-holomorphic curve is an associative submanifold, whereas the cone of a Lagrangian submanifold is a coassiciative submanifold.
@@ 6.1. Totally geodesic submanifolds of Riemannian cones
We now discuss the relationship between the totally geodesic submanifolds of a Riemannian cone (over a complete real analytic manifold) and those of its base. We will be interested in determining the maximal totally geodesic submanifolds of the cone $\widehat{M}$ over $M$.
@@
It was shown in [32] that certain types of Riemannian manifolds with special holonomy do not admit totally geodesic hypersurfaces. We deduce that the same result holds for Sasakian-Einstein, $6$-dimensional nearly Kähler and nearly parallel $\mathsf{G}_{2}$-manifolds.
@@
    $M^{6}$ is a $6$-dimensional strictly nearly Kähler manifold,
@@
Let $\Sigma$ be a totally geodesic hypersurface of $M$. By Lemma 6.6, $\widehat{\Sigma}$ is a totally geodesic hypersurface of $\widehat{M}$. By the Gallot Theorem (see [20]), $\widehat{M}$ is locally irreducible since $M$ has non-constant sectional curvature. Moreover, by [5], we know that:
@@
    If $M^{2n+1}$ is a Sasakian-Einstein manifold, then the restricted holonomy of $\widehat{M}$ is contained in $\mathsf{SU}(n+1)$.
@@
Now, since $\widehat{M}$ is Einstein (indeed Ricci-flat), by [32, Theorem 4.3] the restricted holonomy of the cone $\widehat{M}$ is $\mathsf{SO}(T_{p}\widehat{M})$. This contradicts the fact that the holonomy of $\widehat{M}$ is contained in one of three aforementioned groups, yielding the result. ∎
@@
Secondly, assume that $(\partial_{r})\notin V$, so that $V$ is a hyperplane in $\mathbb{R}(\partial_{r})_{(\tau,p)}\oplus W$. In this setting, $\Sigma$ is complete by Corollary 3.5 and for all $y\in\Sigma$ the tangent space $\widetilde{f}(y)$ does not contain the vector $(\partial_{r})_{f(y)}$ (otherwise, $\Sigma$ would admit a noncomplete geodesic). Thus, the radial vector field is nowhere tangent to $f\colon\Sigma\to M$. We now argue in a similar way as in the previous paragraph. Let $\varepsilon>0$ be such that $\widehat{\exp}_{(\tau,p)}$ is a diffeomorphism of $B_{T_{(\tau,p)}\widehat{M}}(0,\varepsilon)$ to its image and $F=\widehat{\exp}_{(\tau,p)}(V\cap B_{T_{(\tau,p)}\widehat{M}}(0,\varepsilon))$ is an embedded totally geodesic submanifold of $\widehat{M}$. Then, as $\partial_{r}$ is nowhere tangent to $F$, the restriction of $\pi$ to $M$ has constant rank equal to $k$, so we may shrink $\varepsilon$ so as to have that $\pi(F)$ is a $k$-dimensional embedded submanifold of $M$ and $\pi\colon F\to\pi(F)$ is a diffeomorphism. The same argument as above shows that $\pi(F)$ is a totally geodesic submanifold, so in particular $W$ is a totally geodesic subspace of $T_{p}M$. Let $g\colon S\to M$ be the associated complete compatible totally geodesic extension and $y\in S$ the unique point with $g(y)=(\tau,p)$ and $\widetilde{g}(y)=W$. Then the cone map $\widehat{g}\colon\widehat{S}\to\widehat{M}$ is the inextendable compatible totally geodesic immersion associated with $\mathbb{R}(\partial_{r})_{(\tau,p)}\oplus W$. Because $V\subseteq\mathbb{R}(\partial_{r})_{(\tau,p)}\oplus W$, we may use Proposition 3.6 to conclude. ∎
@@
Theorem 6.9 reduces the classification of (maximal) totally geodesic submanifolds of cones to that of the totally geodesic submanifolds of the base manifold, and separately to that of totally geodesic hypersurfaces in the cone. We note that these hypersurfaces may not arise as cones over totally geodesic hypersurfaces in the base space, as we will see in Example 6.13 and Example 6.15.
@@
Due to Theorem 6.9, we may focus only on totally geodesic hypersurfaces. Let $M$ be a connected complete Riemannian manifold of constant sectional curvature $\kappa\in\mathbb{R}$ and dimension $n\geq 2$, and suppose that $\widehat{M}$ admits a totally geodesic hypersurface $\Sigma$ that is not tangent to the cone vector field. By shrinking $\Sigma$ if necessary, we can assume that $\Sigma\subseteq\widehat{M}$ is embedded and every $(\tau,p)\in\Sigma$ is such that $V=T_{(\tau,p)}\Sigma$ does not contain $\partial_{r}$, so its orthogonal complement must be generated by a vector of the form $\partial_{r}+\eta$ for a certain $\eta\in T_{p}M$. Now, from Remark 6.11 and (9) we deduce that $0=\widehat{R}(\eta,v)v=(\kappa-1)(\lvert v\rvert^{2}\eta-\langle\eta,v\rangle v)$ for all $v\in T_{p}M$, which means that either $\kappa=1$ or $\eta=0$. If $\kappa\neq 1$, we deduce that $V=T_{(\tau,p)}M$ for all $(\tau,p)\in\Sigma$, so $\Sigma$ is an integral manifold of the distribution $\mathcal{D}=\partial_{r}^{\perp}$ on $\widehat{M}$. The maximal integral manifolds of $\mathcal{D}$ are precisely the links $\{\tau\}\times M$ for each $\tau>0$, so $\Sigma$ is actually an open subset of a leaf $\{\tau_{0}\}\times M$ for a certain $\tau_{0}>0$. However, the last equation in (8) shows that the leaves are never totally geodesic, so we arrive at a contradiction. We conclude that if $\kappa\neq 1$, the inextendable totally geodesic hypersurfaces of $\widehat{M}$ are precisely the cones over the complete totally geodesic hypersurfaces of $M$.
@@
Let $M$ be equal to either $\mathsf{S}^{3}_{\mathbb{C},\tau}(r)$ or $\mathbb{R}\mathsf{P}^{3}_{\mathbb{C},\tau}(r)$ and let $\widehat{M}$ denote the Riemannian cone over $M$. Then, $\widehat{M}$ admits a totally geodesic hypersurface if and only if $\tau=r=1$.
@@
Let $M=\mathsf{S}^{3}_{\mathbb{C},\tau}(r)$ be a three-dimensional Berger sphere of radius $r$ and deformation parameter $\tau$. We will show that $\widehat{M}$ does not admit totally geodesic hypersurfaces unless $r=\tau=1$ (that is, $M$ is the unit round sphere).
@@
Let us suppose that $M\neq\mathsf{S}^{3}(1)$ and $\Sigma$ is a totally geodesic hypersurface of $\widehat{M}$. We may assume that $\Sigma$ is embedded in $M$. Because $M$ is homogeneous, Corollary 6.4 allows us to suppose that $\Sigma$ passes through a point of the form $(t,o)$ with tangent space $V\subseteq T_{(t,o)}\widehat{M}\equiv\mathbb{R}\partial_{r}\oplus\mathfrak{p}$. As $M$ does not admit totally geodesic hypersurfaces [48, Theorem A], we have that $\partial_{r}\notin V$, so $V^{\perp}$ must be spanned by a vector of the form $\partial_{r}+\eta$, where $\eta\in\mathfrak{p}$. We may write $\eta=a_{1}E+a_{2}X+a_{3}Y$ for some constants $a_{1}$, $a_{2}$, $a_{3}\in\mathbb{R}$. From Remark 6.11 we also know that $\widehat{R}(\eta,Z)Z=0$ for all $Z\in\mathfrak{p}$. A polarization argument shows that the previous condition is equivalent to $\widehat{R}(\eta,Z)W+\widehat{R}(\eta,W)Z=0$ for all $Z$, $W\in\mathfrak{p}$.
@@
In view of these latter results, one could think that these totally geodesic hypersurfaces nowhere tangent to the radial vector can only happen if the base is a sphere of radius one. However, the following example shows that there are many “non-trivial” local examples of such phenomena.

### v3

@@ 6. Riemannian cones and totally geodesic submanifolds
In this section we start by recalling the definition and the basic properties of Riemannian cones. After that we prove a structure result for totally geodesic submanifolds of Riemannian cones. We refer the reader to [41] for a detailed account on semi-Riemannian cones.
Let $M$ be a Riemannian manifold, which for our purposes is assumed to be real analytic and complete. We define its Riemannian cone as the warped product $\widehat{M}=\mathbb{R}^{+}\times_{f}M$, where $f\colon\mathbb{R}^{+}\to\mathbb{R}^{+}$ is the identity map. More explicitly, if $g=\langle\cdot,\cdot\rangle$ denotes the metric on $M$ and $r\colon\widehat{M}\to\mathbb{R}^{+}$ is the projection on the first factor, the inner product on $\widehat{M}$ is given by $\widehat{g}=\mathrm{d}r^{2}+r^{2}g$.
@@
We can also describe the geodesics of $\widehat{M}$ in terms of those of $M$. Let $(\tau,p)\in\widehat{M}$ be any point and consider the tangent vector $w=a\partial_{r}+v$, where $a\in\mathbb{R}$ and $v\in T_{p}M$ are arbitrary. From [41, Equation 2.7], we see that the geodesic $\widehat{\gamma}(t)=\widehat{\exp}_{(\tau,p)}(tw)=(\rho(t),\alpha(t))$ is given in a neighborhood of $t=0$ by
@@
Now, let $f\colon\widehat{M}\to\widehat{N}$ be an isometry and fix $(\tau,p)\in\widehat{M}$ with image $(s,q)=f(\tau,p)$. Since $f$ is an isometry, it sends $\mathcal{C}_{(\tau,p)}$ to $\mathcal{C}_{(s,q)}$, and thus we have $f_{*(\tau,p)}((\partial_{r})_{(\tau,p)})=\pm(\partial_{r})_{(s,q)}$. The equality $f_{*(\tau,p)}((\partial_{r})_{(\tau,p)})=~{}-(\partial_{r})_{(s,q)}$ is not possible, because in that case the maximal geodesic $\gamma(t)=\widehat{\operatorname{exp}}_{(\tau,p)}(t\partial_{r})$ would be mapped to the maximal geodesic $\beta(t)=\widehat{\operatorname{exp}}_{(s,q)}(-t\partial_{r})$, which is not possible because the first geodesic is defined for all $t\geq 0$, while the second one is not. We deduce that $f_{*(\tau,p)}((\partial_{r})_{(\tau,p)})=~{}(\partial_{r})_{f(\tau,p)}$. In addition, the maximal geodesic $\gamma(t)=\widehat{\operatorname{exp}}_{(\tau,p)}(t\partial_{r})$ is mapped to the maximal geodesic $\beta(t)=~{}\widehat{\operatorname{exp}}_{(s,q)}(t\partial_{r})$, and their corresponding intervals of definition are $(-\tau,0)$ and $(-s,0)$. Hence, $\tau=s$. All in all, we have seen that $f$ sends the link $\{\tau\}\times M$ to $\{\tau\}\times N$ for each $\tau\in\mathbb{R}^{+}$, so $f$ takes the form $f(\tau,p)=(\tau,h(\tau,p))$ for a map $h\colon\widehat{M}\to N$. Furthermore, at each $(\tau,p)\in\widehat{M}$ we have
@@
Many geometric properties of Riemannian manifolds can be translated into geometric properties of their cones. For instance, as a consequence of (9), $M$ is an Einstein manifold with $\operatorname{Ric}=(n-1)\langle\cdot,\cdot\rangle$ if and only if $\widehat{M}$ is Ricci-flat. It turns out that nearly Kähler structures on six-dimensional manifolds are related to $\mathsf{G}_{2}$-structures on their cones. We briefly describe this relationship, see [5] for details. For the general theory of $\mathsf{G}_{2}$-manifolds, we refer the reader to [32, Chapter 10].
@@
There is also a relationship between submanifolds of $M$ that have a nice interaction with $J$ and calibrated cones inside $\widehat{M}$. The notions of calibrated geometry were introduced in the seminal paper [26] by Harvey and Lawson. We remind that a calibration on a Riemannian manifold $N$ is a closed differential form $\omega\in\Omega^{k}(M)$ satisfying $\omega(v_{1},\dots,v_{k})\leq 1$ whenever $v_{1},\dots,v_{k}$ are unit vectors in $TN$. A $k$-dimensional oriented submanifold $S$ of $N$ is calibrated if the restriction of $\omega$ to $S$ is equal to the Riemannian volume form of $S$, and it follows that $S$ is a minimal submanifold. It can be shown that for the case of a $\mathsf{G}_{2}$-manifold $(N,\phi)$, both $\phi$ and its Hodge dual $\star\phi$ are calibrations ([26, Theorem 1.4 and Theorem 1.16]). We say in this case that $S$ is associative (respectively, coassociative) if it is calibrated with respect to $\phi$ (respectively, $\star\phi$). Coming back to the case that $N=\widehat{M}$ is the cone of a six-dimensional strict nearly Kähler manifold, it is known that the cone of a $J$-holomorphic curve is an associative submanifold, whereas the cone of a Lagrangian submanifold is a coassociative submanifold.
@@ 6.1. Totally geodesic submanifolds of Riemannian cones
We now discuss the relationship between the totally geodesic submanifolds of a Riemannian cone (over a complete real analytic manifold) and those of its base. We are interested in determining the maximal totally geodesic submanifolds of the cone $\widehat{M}$ over $M$.
@@
It was shown in [31] that certain types of Riemannian manifolds with special holonomy do not admit totally geodesic hypersurfaces. We deduce that the same result holds for Sasakian–Einstein, $6$-dimensional nearly Kähler and nearly parallel $\mathsf{G}_{2}$-manifolds.
@@
    $M^{6}$ is a $6$-dimensional strictly nearly Kähler manifold, or
@@
Let $\Sigma$ be a totally geodesic hypersurface of $M$. By Lemma 6.6, $\widehat{\Sigma}$ is a totally geodesic hypersurface of $\widehat{M}$. By Gallot’s theorem (see [20]), $\widehat{M}$ is locally irreducible since $M$ has non-constant sectional curvature. Moreover, by [5], we know that:
@@
    If $M^{2n+1}$ is a Sasakian–Einstein manifold, then the restricted holonomy of $\widehat{M}$ is contained in $\mathsf{SU}(n+1)$.
@@
Now, since $\widehat{M}$ is Einstein (indeed Ricci-flat), by [31, Theorem 4.3] the restricted holonomy of the cone $\widehat{M}$ is $\mathsf{SO}(T_{p}\widehat{M})$. This contradicts the fact that the holonomy of $\widehat{M}$ is contained in one of three aforementioned groups, yielding the result. ∎
@@
Secondly, assume that $(\partial_{r})_{(\tau,p)}\notin V$, so that $V$ is a hyperplane in $\mathbb{R}(\partial_{r})_{(\tau,p)}\oplus W$. In this setting, $\Sigma$ is complete by Corollary 3.5 and for all $y\in\Sigma$ the tangent space $\widetilde{f}(y)$ does not contain the vector $(\partial_{r})_{f(y)}$ (otherwise, $\Sigma$ would admit a noncomplete geodesic). Thus, the radial vector field is nowhere tangent to $f\colon\Sigma\to M$. We now argue in a similar way as in the previous paragraph. Let $\varepsilon>0$ be such that $\widehat{\exp}_{(\tau,p)}$ is a diffeomorphism of $B_{T_{(\tau,p)}\widehat{M}}(0,\varepsilon)$ to its image and $F=\widehat{\exp}_{(\tau,p)}(V\cap B_{T_{(\tau,p)}\widehat{M}}(0,\varepsilon))$ is an embedded totally geodesic submanifold of $\widehat{M}$. Then, as $\partial_{r}$ is nowhere tangent to $F$, the restriction of $\pi$ to $M$ has constant rank equal to $k$, so we may shrink $\varepsilon$ so as to have that $\pi(F)$ is a $k$-dimensional embedded submanifold of $M$ and $\pi\colon F\to\pi(F)$ is a diffeomorphism. The same argument as above shows that $\pi(F)$ is a totally geodesic submanifold, so in particular $W$ is a totally geodesic subspace of $T_{p}M$. Let $g\colon S\to M$ be the associated complete compatible totally geodesic extension and $y\in S$ the unique point with $g(y)=(\tau,p)$ and $\widetilde{g}(y)=W$. Then the cone map $\widehat{g}\colon\widehat{S}\to\widehat{M}$ is the inextendable compatible totally geodesic immersion associated with $\mathbb{R}(\partial_{r})_{(\tau,p)}\oplus W$. Because $V\subseteq\mathbb{R}(\partial_{r})_{(\tau,p)}\oplus W$, we may use Proposition 3.6 to conclude. ∎
@@
Theorem 6.9 reduces the classification of (maximal) totally geodesic submanifolds of cones to that of totally geodesic submanifolds of the base manifold, and separately to that of totally geodesic hypersurfaces. We note that these hypersurfaces may not arise as cones over totally geodesic hypersurfaces in the base space, as we will see in Example 6.13 and Example 6.15.
@@
Due to Theorem 6.9, we may focus only on totally geodesic hypersurfaces. Let $M$ be a connected complete Riemannian manifold of constant sectional curvature $\kappa\in\mathbb{R}$ and dimension $n\geq 2$, and suppose that $\widehat{M}$ admits a totally geodesic hypersurface $\Sigma$ that is not tangent to the cone vector field. By shrinking $\Sigma$ if necessary, we can assume that $\Sigma\subseteq\widehat{M}$ is embedded and every $(\tau,p)\in\Sigma$ is such that $V=T_{(\tau,p)}\Sigma$ does not contain $\partial_{r}$, so its orthogonal complement must be generated by a vector of the form $\partial_{r}+\eta$ for a certain $\eta\in T_{p}M$. Now, from Remark 6.11 and (9) we deduce that $0=\widehat{R}(\eta,v)v=(\kappa-1)(\lvert v\rvert^{2}\eta-\langle\eta,v\rangle v)$ for all $v\in T_{p}M$, which means that either $\kappa=1$ or $\eta=0$. If $\kappa\neq 1$, we obtain that $V=T_{(\tau,p)}M$ for all $(\tau,p)\in\Sigma$, so $\Sigma$ is an integral manifold of the distribution $\mathcal{D}=\partial_{r}^{\perp}$ on $\widehat{M}$. The maximal integral manifolds of $\mathcal{D}$ are precisely the links $\{\tau\}\times M$ for each $\tau>0$, so $\Sigma$ is actually an open subset of a leaf $\{\tau_{0}\}\times M$ for a certain $\tau_{0}>0$. However, the last equation in (8) shows that the leaves are never totally geodesic, so we arrive at a contradiction. We conclude that if $\kappa\neq 1$, the inextendable totally geodesic hypersurfaces of $\widehat{M}$ are precisely the cones over the complete totally geodesic hypersurfaces of $M$.
@@
Let $M$ be either $\mathsf{S}^{3}_{\mathbb{C},\tau}(r)$ or $\mathbb{R}\mathsf{P}^{3}_{\mathbb{C},\tau}(r)$ and let $\widehat{M}$ denote the Riemannian cone over $M$. Then, $\widehat{M}$ admits a totally geodesic hypersurface if and only if $\tau=r=1$.
@@
Let $M=\mathsf{S}^{3}_{\mathbb{C},\tau}(r)$ be a three-dimensional Berger sphere of radius $r$ and deformation parameter $\tau$. We show that $\widehat{M}$ does not admit totally geodesic hypersurfaces unless $r=\tau=1$ (that is, $M$ is the unit round sphere).
@@
Let us suppose that $M\neq\mathsf{S}^{3}(1)$ and $\Sigma$ is a totally geodesic hypersurface of $\widehat{M}$. We may assume that $\Sigma$ is embedded in $M$. Because $M$ is homogeneous, Corollary 6.4 allows us to suppose that $\Sigma$ passes through a point of the form $(t,o)$ with tangent space $V\subseteq T_{(t,o)}\widehat{M}\equiv\mathbb{R}\partial_{r}\oplus\mathfrak{p}$. As $M$ does not admit totally geodesic hypersurfaces [46, Theorem A], we have that $\partial_{r}\notin V$, so $V^{\perp}$ must be spanned by a vector of the form $\partial_{r}+\eta$, where $\eta\in\mathfrak{p}$. We may write $\eta=a_{1}E+a_{2}X+a_{3}Y$ for some constants $a_{1}$, $a_{2}$, $a_{3}\in\mathbb{R}$. From Remark 6.11 we also know that $\widehat{R}(\eta,Z)Z=0$ for all $Z\in\mathfrak{p}$. A polarization argument shows that the previous condition is equivalent to $\widehat{R}(\eta,Z)W+\widehat{R}(\eta,W)Z=0$ for all $Z$, $W\in\mathfrak{p}$.
@@
In view of the latter results, one could think that totally geodesic hypersurfaces nowhere tangent to the radial vector can only appear if the base is a sphere of radius one. However, the following construction shows that there are many nontrivial local examples of totally geodesic hypersurfaces satisfying this condition.